OpenStudy (anonymous):

Solve SA = 2ab + 2bc + 2ac for b

OpenStudy (unklerhaukus):

factor out the b from the first two terms

OpenStudy (anonymous):

how?

OpenStudy (unklerhaukus):

SA = 2ab + 2bc + 2ac =b(2a+2c)+2ac

OpenStudy (unklerhaukus):

now take away that last term from both sides, and finally divide both sides of the equation by the factor that leaves only b on the RHS

OpenStudy (anonymous):

-2ac=b(2ac+2c?

OpenStudy (anonymous):

what does SA stand for?

jimthompson5910 (jim_thompson5910):

It should look something like this: \[\Large SA = 2ab + 2bc + 2ac\] \[\Large SA = b(2a + 2c) + 2ac\] \[\Large SA-2ac = b(2a + 2c)\] \[\Large b(2a + 2c) = SA-2ac\] What's the last step?

OpenStudy (anonymous):

it's part of the problem.

OpenStudy (anonymous):

divide by -2ac?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

you have to divide both sides by 2a+2c

jimthompson5910 (jim_thompson5910):

So here are the full/complete steps to isolating b: \[\Large SA = 2ab + 2bc + 2ac\] \[\Large SA = b(2a + 2c) + 2ac\] \[\Large SA-2ac = b(2a + 2c)\] \[\Large b(2a + 2c) = SA-2ac\] \[\Large b = \frac{SA-2ac}{2a + 2c}\]

OpenStudy (anonymous):

that doesnt match any of my answer choices?

jimthompson5910 (jim_thompson5910):

What are your answer choices

OpenStudy (anonymous):

a.)SA/2(a+c) b.)SA-4a/c c.)1/2 Sa-ac/a+c d.) 1/2SA/a+b+c

jimthompson5910 (jim_thompson5910):

oh, for some reason, they're factoring out 2's (not sure why) \[\Large b = \frac{SA-2ac}{2a + 2c}\] \[\Large b = \frac{2(\frac{1}{2}SA-ac)}{2(a + c)}\] \[\Large b = \frac{\frac{1}{2}SA-ac}{a + c}\]

OpenStudy (anonymous):

thank you!

OpenStudy (unklerhaukus):

@BlahblerBlah, SA stands for the surface area of the prism